A hydrostatic bearing is used to rotate a load. The load 500kN, fluid viscosity 500 mPa-s. The bearing has an outside diameter of 500mm and an inside diameter of 400mm.

a) What flow Q is required if the film thickness is 1mm?
b) What Pressure is required?
c) what is the friction torque if it is rotated at 3 rad/s?
1
Expert's answer
2019-11-18T08:25:23-0500
"Fout=\pi*D^2/4=3.1415*500/4=196343.75"

"Fin=\pi*d^2/4=3.1415*400^2/4=125660"

"\Delta F=Fout-Fin=70683.75"


"Feff=Fin+0.5*\Delta F=125660+0.5*70683.75=161002"

PRESSURE

"p=N\/Feff=500000N\/161002mm^2=3.11MPa"

FLOW

"Q=(p*h^3/(12*\rho))*((D-d)/2)*(1/(\pi*D))=3.11MPa*(1mm)^3*(500mm-400mm)/(12*500*10^(-9) MPa*s*2*3.1415*500mm)=16500mm^3/s=0.0165l/s"

FRICTION TORQUE

"M=\rho*\Delta F*\omega*(D-d)^2/(h*2*2^2)=500*10^(-9)MPa*s*70684mm^2*3rad/s*(500mm-400mm)^2/(1mm*2*4)=10800N*mm=10.8Nm"